# SoccerDiffusion: Toward Learning End-to-End Humanoid Robot Soccer from Gameplay Recordings

### v1

@@ 4.2 Qualitative Evaluation
For the qualitative evaluation, we used one Wolfgang-OP robot in the real world. The robot was able to walk in a stable manner and perform a variety of motions, such as kicking, turning, standing up, and walking in all directions. It showed dynamic motions and stabilization behaviors mimicking the way the Hamburg Bit-Bots stabilize their robot (PID-based Center of Mass control and adaptive gait phase) and is able to recover from disturbances. However, it is hard to quantify the performance of the robot in this case due to the end-to-end nature of the model: Evaluating the walking speed or the traversed distance is not meaningful because we can not influence the model, e.g., to walk with maximum speed, and we do not know the intended behavior.
The robots motions are very similar to the ones in the dataset, but the robot fails to show any high level behavior, such as intentional kicking or positioning itself on the soccer field. Possible reasons could be a limited number of demonstrations, the model size, or the data distribution itself. Due to team play, not all players are moving towards the ball at all times, which increases the complexity of the dataset and might be hard to learn for the model.
The image sequence in Figure 2 shows the robot walking forward, while Figure 3 shows the robot’s recovery from a fall to the side, both using our SoccerDiffusion policy. When executing the falling and standing up motion the robot performs a complex multi-step motion (i. e. turning itself from the side to the front and then standing up).
@@
Figure 2: Bipedal walk cycle, on the Wolfgang-OP platform [2].
Figure 3: Fall recovery from the right side, on the Wolfgang-OP platform [2].
Different head motions imitating the training data can also be observed. Firstly, the robot performs a head-scan motion to the left and right at various tilt angles. This is usually done to scan the environment for more accurate localization. Secondly, the robot performs a tilting behavior of the head to observe the ball and the environment. This behavior is mimicking the head behavior during ball manipulation in the dataset. The robot does not follow the ball with its head. This is to be expected since the demonstration does not track the ball either.
A major limitation of the current model is a strong bias to keep standing still, if there was no prior motion in the context window. This is likely due to the fact that the dataset contains a lot of standing robots at the start of the game or in game for e.g. the goalkeeper who stands in the goal most of the time. In combination with lack of high level behavior, this leads to the robot rarely starting to walk on its own.
Overall, the learned motions seem to be very robust, as the model is able to overcome the gap between the real world and simulation without any additional training on simulated data (see Table 2). The software stack of the Hamburg Bit-Bots interestingly does not exhibit this behavior, as the walk engine parameter needs to be tuned for the simulation.
@@ 5 Future Work
Since we could not reach the same level of performance as the Hamburg Bit-Bots, we recognize the significant potential for future improvements. Currently, we are working on integrating large-scale datasets from other teams. Furthermore, employing a general-purpose vision embedding model such as DINOv2 [14] could facilitate the learning of more complex, high-level behaviors while also reducing the computational demands of training, as the embeddings can be precomputed. Additionally, we could filter the dataset, ignoring unnecessary data, such as the robot being stopped or the ball being out of play. While this is not the primary focus of this work, auxiliary losses for predicting intermediate representations, such as the robot’s position or walk velocity, could be beneficial. As previously mentioned, the model seems well suited for follow-up work using reinforcement learning to further improve performance.
@@ 6 Conclusion
Our model successfully learned to reproduce a range of fundamental behaviors observed in the dataset. The generated motions included stable walking in various directions, turning, kicking motions, and robust fall recovery sequences. Notably, the model was able to replicate stabilization behaviors to maintain balance during walking. Furthermore, we show real-to-sim transfer of the learned model. Through the use of distillation, we were able to reduce computational cost to achieve real-time performance.
We laid a foundation for further optimization of an end-to-end policy for humanoid robot soccer using RL or PO techniques. However, we did not achieve the same level of behavior as the Hamburg Bit-Bots, such as intentional kicking or positioning on the field. We suspect this is due to the limited dataset and will be addressed in future work.
@@ 6.0.1 Acknowledgements
We gratefully acknowledge funding and support from the project Digital and Data Literacy in Teaching Lab (DDLitLab) at the University of Hamburg and the Stiftung Innovation in der Hochschullehre foundation. We extend our special thanks to the members of the Hamburg Bit-Bots RoboCup team for their continuous support and for providing data and computational resources. We also thank the RoboCup teams B-Human and HULKs for generously sharing their data for this research. Additionally, we are grateful to the Technical Aspects of Multimodal Systems (TAMS) research group at the University of Hamburg for providing computational resources. This research was partially funded by the Ministry of Science, Research and Equalities of Hamburg, as well as the German Research Foundation (DFG) and the National Science Foundation of China (NSFC) through the project Crossmodal Learning (TRR-169).

### v2

@@ 4.2 Qualitative Evaluation
For the qualitative evaluation, we used one Wolfgang-OP robot in the real world. The robot was able to walk in a stable manner and perform various motions, such as kicking, turning, standing up, and walking in all directions. It showed dynamic motions and stabilization behaviors mimicking the way the Hamburg Bit-Bots stabilize their robot (PID-based Center of Mass control and adaptive gait phase) and is able to recover from disturbances. However, it is difficult to quantify the performance for these motions due to the end-to-end nature of the model: Evaluating the walking speed or the traversed distance is not meaningful because we cannot influence the model, e.g., to walk at maximum speed, and we do not know the intended behavior.
The robot’s motions are very similar to the ones in the dataset, but the robot fails to show any high-level behavior, such as intentional kicking or positioning itself on the soccer field. Possible reasons could be a limited number of demonstrations, the model size, or the data distribution itself. Due to team play, not all players are moving towards the ball at all times, which increases the complexity of the dataset and might be difficult to learn for the model.
The image sequence in Fig. 2 shows the robot walking forward and recovering from a fall to the side, both using our SoccerDiffusion policy. When executing the falling and standing up motion, the robot performs a multistep motion (i.e., turning itself from the side to the front and then standing up).
@@
(a) Bipedal walk cycle
(b) Fall recovery from the right side
Figure 2: Qualitative evaluation: (a) walking and (b) fall recovery, both performed by the SoccerDiffusion policy on the Wolfgang-OP platform [2].
Different head motions imitating the training data can be observed. Firstly, the robot performs a head-scan motion to the left and right side at various tilt angles. This is usually done to scan the environment for more accurate localization. Secondly, the robot tilts the head to observe the ball and the environment. This behavior is mimicking the head behavior during ball manipulation in the dataset. The robot does not follow the ball with its head, which is to be expected since robots in the training data do not track the ball either.
A major limitation of the current model is a strong bias to keep standing still, if there was no prior motion in the context window. This is likely because the dataset contains many sequences of standing robots. Together with a lack of high-level behavior, this leads to the robot rarely starting to walk on its own.
Overall, the learned motions are seemingly very robust, as the model can overcome the gap between the real world and simulation without any additional training on simulated data (see Table 3). The software stack of the Hamburg Bit-Bots interestingly does not exhibit this behavior, as the walk engine parameters need to be tuned for the simulation.
@@ 5 Future Work
Since we could not reach the same level of abilities as the conventional baseline, we recognize significant potential for future improvements. Currently, we are expanding our dataset with recordings from other teams and 3D-simulated autonomous matches. Furthermore, employing a general-purpose vision embedding model such as DINOv2 [17] could facilitate the learning of more complex, high-level behaviors while also reducing the computational demands of training, as the embeddings can be precomputed. From experience, we know the dataset contains long sequences of a robot standing still during the game. We could filter out these sequences to encourage more dynamic actions. While this is not the primary focus of this work, auxiliary losses for predicting intermediate representations, such as the robot’s position or walk velocity, could be beneficial. A detailed ablation study could identify more effective configurations for our task.
@@ 6 Conclusion
Our model successfully learned to reproduce a range of fundamental behaviors from the dataset. The generated motions included stable walking in various directions, turning, kicking motions, and robust fall recovery sequences. Notably, the model replicates stabilization behaviors to maintain balance while walking. Furthermore, we show real-to-sim transfer of the learned model. We achieved real-time performance by lowering computational costs through distillation.
We laid a foundation for further optimization of an end-to-end policy for humanoid robot soccer using RL or PO techniques. However, we did not achieve the same level of behavior as the Hamburg Bit-Bots, such as intentional kicking or positioning on the field. We suspect this is due to the limited dataset or training duration and will be addressed in future work.
@@ 6.0.1 Acknowledgements
We gratefully acknowledge funding and support from the project Digital and Data Literacy in Teaching Lab (DDLitLab) at the University of Hamburg and the Stiftung Innovation in der Hochschullehre foundation. We extend our special thanks to the members of the Hamburg Bit-Bots RoboCup team for their continuous support and for providing data and computational resources. We also thank the RoboCup teams B-Human and HULKs for generously sharing their data for this research. Additionally, we are grateful to the Technical Aspects of Multimodal Systems (TAMS) research group at the University of Hamburg for providing computational resources. This research was partially funded by the Ministry of Science, Research and Equalities of Hamburg, as well as the German Research Foundation and the National Science Foundation of China through the project Crossmodal Learning (TRR-169).